OpenStudy (anonymous):

can someone describe to me how to find the derivative of an inverse function?

OpenStudy (anonymous):

Question?

OpenStudy (anonymous):

Question?

OpenStudy (anonymous):

Follow this...

OpenStudy (anonymous):

suppose \[y = f(x) = 2x ^{2} - 3x\] if h(x) is the inverse function of f, then h'(-1) equals...

OpenStudy (jamesj):

In general, the derivative of the inverse function is \[ (f^{-1})'(a) = \frac{1}{f'(f^{-1}(a)} \] Can you do it now?

OpenStudy (anonymous):

i don't understand how to do \[f(x ^{-1})\]

OpenStudy (jamesj):

In your notation \[ f(x) = 2x^2 - 3x \] and h is the inverse, \[ h = f^{-1} \] Hence \[ h'(-1) = \frac{1}{f'(h(1))} \] So you first step is to find h(1). I.e., for what value(s) of x is f(x) = 1?

OpenStudy (anonymous):

oh! okay

OpenStudy (mathteacher1729):

mvkellogg -- do you understand the geometry of this concept? (how the derivative of a function and its inverse are related)

OpenStudy (jamesj):

(there should be a minus sign in my last equation; I'm sure you figured that out already.)

OpenStudy (anonymous):

mathteacher1729.. kind of....

OpenStudy (mathteacher1729):

Long story short: f(x) has a tangent line of slope m at the point (x,y). f inverse has a tangent line of slope (-1/m) at point (y,x). I'm going to try to find a good pic of this somewhere.

OpenStudy (anonymous):

alright